Assume that individuals have a 3% chance of suffering from a specific allergy.

a. Out of a population of 100, what is the chance that fewer than 2 people suffer from the allergy?

b. Approximate the number of allergy sufferers in part a) using the Poisson distribution.
Answer #1

TOPIC:Application of Binomial and Poisson distribution.

- ax Define, p = P (of suffering from specific allergy) = 0.03 Here, n = sample size = 100. Ket, X = No. of people, sufferingso, the sound of yis- > 720,1,2,3,... ( 0 jotherwise, . We need, => P (742), = P (YLD. =P(y=0) +8 (7=1). = ē3 + 3xê 3. = 423